### v1

# Bayesian competing risks survival modeling for assessing the cause of death of patients with heart failure

Abstract
Competing risk models are survival models with several events of interest acting in competition and whose occurrence is only observed for the event that occurs first in time. This paper presents a Bayesian approach to these models in which the issue of model selection is treated in a special way by proposing generalizations of some of the Bayesian procedures used in univariate survival analysis. This research is motivated by a study on the survival of patients with hearth failure undergoing cardiac resynchronization therapy, a procedure which involves the implant of a device to stabilize the heartbeat. Two different causes of causes of death have been considered: cardiovascular and non-cardiovascular, and a set of baseline covariates are examined in order to better understand their relationship with both causes of death. Model selection procedures and model checking analyses have been implemented and assessed. The posterior distribution of some relevant outputs such as the overall survival function, cumulative incidence functions, and transition probabilities have been computed and discussed.
keywords:
Cardioscular and non-cardiovascular death; Conditional Predictive Ordinate; Model uncertainty; Transition probabilities.
1 Introduction
Cardiac resynchronization therapy (CRT) is an electrical intervention for heart failure patients. This treatment significantly enhances their quality of life, functional capacity, and reduces heart failure-related hospitalizations and cardiovascular mortality. It is particularly indicated in patients with a left ventricular ejection fraction (LVEF) less than or equal to 35% and complex QRS width higher than 130 milliseconds despite optimized pharmacological therapy (Bristow et al., 2004; Cleland et al., 2005; Moss et al., 2009).
LVEF and left ventricular end-diastolic and end-systolic volumes (LVEDV and LVESV, respectively) are commonly used as clinical markers reflecting global left ventricular systolic performance or remodeling. CRT fosters reverse remodeling of the left ventricle, leading to decreased LVESV and improved LVEF. Clinical studies have established that an increase above 5% in LVEF or a reduction of 15% in LVESV are associated with enhanced survival (Cleland et al., 2013; Ghio et al., 2009). These benchmarks are commonly used to define a favorable response to CRT. Nevertheless, despite advances in technology and patient selection, around 30% of patients are classified as non-responders to this therapy, a definition that might now be under debate owing to the observed benefits in lesser degrees of left ventricular remodeling compared to the natural course of the disease.
Different studies agree on reporting a mortality risk of approximately 10% among patients receiving CRT, with heart failure being the leading cause (Marijon et al., 2015; Alvarez-Alvarez et al., 2021). In addition, some researchers have identified male gender, advanced age, ischemic etiology, and severity of mitral regurgitation as indicators of mortality in this cohort (Cleland et al., 2005). However, the risk of mortality and heart failure-related hospitalizations fluctuates during follow-up. Assessing left ventricular (LV) reverse remodeling through longitudinal measurements of LV structure and function over time in heart failure patients presents a more comprehensive method for stratifying cardiovascular mortality and heart failure decompensation risk (Alvarez-Alvarez et al., 2021).
The objective of this paper is to investigate baseline predictors of cardiovascular and non-cardiovascular mortality as well as to assess the survival times in heart failure patients undergoing CRT, knowing that their death may eventually be due to cardiovascular or non-cardiovascular causes.
The statistical framework for the aforementioned objectives is Survival Analysis. In particular, competing risk models in which the event of interest, in our case death, can occur from a variety of competing causes. The statistical literature is abundant regarding procedures for modeling competing survival scenarios (Putter et al., 2007; Chen et al., 2014). In particular, we will focus on cause-specific risk modelling to quantitatively assess the behaviour of the survival process through the cause-specific hazard function for each cause of death, the overall survival function for the time in which death occurs, and the cumulative incidence function of that time due to each of the causes considered among others (Chen et al., 2014). We also incorporate a look at the modelling from the point of view of multistate models (Andersen et al., 2001). These are stochastic processes of continuous time and discrete state space representing the different health conditions considered. We discuss the transition probabilities between the different states and specially conditional transition probabilities when it is known that a patient remains alive a certain time after the CRT treatment. We will work within the Bayesian statistical methodology. This allows quantifying all sources of uncertainty associated with a statistical problem in a natural way through probability distributions.
The rest of the paper is organized as follows. Section 2 presents a retrospective follow-up cardiac resynchronization therapy study and a brief description of the subsequent dataset. Section 3 contains general information about the competing risk model, the main elements of the cause-specific risk approach, and the general Bayesian inferential process. Subsection 3.3, devoted to model uncertainty, presents a special variable selection scenario. In this topic, we extend the proposal by García García-Donato et al (2023) in Cox regression models for computing Bayes factors to competing risk models and employ spike-and-slab selection methods. In addition, we also expand the definition of conditional predictive ordinate (CPO) (Ibrahim et al., 2001) to the general competing risk framework which is needed in order to apply the logarithm of the pseudomarginal likelihood as a tool for model assessment. Section 4 returns to the study of CRT therapy and applies the procedures and proposals of Section 3 to the data of this problem. Finally, Section 5 includes some conclusions.
2 The cardiac resynchronization therapy study
[…]
Clinical follow-up was completed in all patients, from the time each individual underwent CRT implantation (time zero) until death, cardiovascular (CV) or non-cardiovascular (N-CV), or until the date of the end of the follow-up. Survival times were observed in the first case and were right-censored in the second. The median and interquartile ranges of survival times were 2.75 and (1.2, 5.2) years, respectively.
Patients were individually followed up in the hospital’s arrhythmia unit to verify the proper functioning of the device. Subsequently, visits were conducted to check the battery status, the patient’s rhythm, stimulation percentage, stimulation and sensing thresholds in each chamber, and electrode impedances. Arrhythmic episodes and corresponding electrograms were reviewed, with different adjustments made based on each patient’s condition. Long-term follow-up included a review of the medical history, which encompassed the development of atrial fibrillation, cardiovascular admissions - including hospitalizations due to heart failure - and mortality, specifying the cause of death. Additionally, the duration of the subsequent QRS on surface electrocardiogram and the evolution of echocardiographic variables were collected through repeated measurements throughout the follow-up period.
A retrospective review was also conducted for each of the patients in the study, including electronic medical reports as well as older records available only in paper format. All medical interventions and hospital records were documented. Vital status was ascertained through telephone calls in the absence of medical reports during follow-up. Furthermore, pharmacological treatment was assessed using the IANUS electronic registry (a database containing patient medical history, including reviews and treatments).
The covariates considered in the study were selected by the team cardiologist. They included age and sex of all patients at the time of implantation, etiology of heart failure (ischemic vs. non-ischemic or dilated), NYHA functional class, creatinine level in blood as a biomarker of renal function, QRS width at the moment of implantation, the type of device which has been implanted (pacemaker or implantable cardioverter-defibrillator ) and the possible occurrence of a left bundle branch block (LBBB) that generates disruptions to the electronic impulse that causes the heart to beat. Additionally, the history of previous atrial fibrillation (Previous AF) has been registered.
Table 1: Descriptive summary of study covariates for censored individuals, for those who died, from CV or N-CV causes, and for all of the together. In the case of quantitative covariates, the sample mean and standard deviation are presented; for categorical variables, the number of patients in the sample in each category and their percentage are reported.
Censored CV Death N-CV Death Total
$n$ 203 55 38 296
Age 68.58 (10.00) 72.27 (7.73) 74.18 (7.15) 69.99 (9.51)
Sex
- Man 150 (73.89%) 48 (87.27%) 29 (76.32%) 227 (76.69%)
- Woman 53 (26.11%) 7 (12.73%) 9 (23.68%) 69 (23.31%)
Creatinine 1.24 (0.45) 1.70 (0.84) 1.39 (0.49) 1.34 (0.57)
CRT Device
- Pacemaker 87 (42.86%) 36 (65.45%) 17 (44.74%) 140 (47.30%)
- Defibrillator 116 (57.14%) 19 (34.55%) 21 (55.26%) 156 (52.70%)
Etiology
- Ischemic 62 (30.54%) 28 (50.91%) 16 (42.11%) 106 (35.81%)
- Dilated 141 (69.46%) 27 (49.09%) 22 (58.89%) 190 (64.19%)
LBBB 80 (39.41%) 26 (47.27%) 12 (31.58%) 118 (39.86%)
NYHA
- Class II 70 (34.48%) 3 (5.45%) 4 (10.53%) 77 (26.01%)
- Class III 126 (62.07%) 43 (78.18%) 33 (86.84%) 202 (68.24%)
- Class IV 7 (3.45%) 9 (16.36%) 1 (2.63%) 17 (5.74%)
Previous AF 83 (40.89%) 31 (56.36%) 24 (63.16%) 138 (46.62%)
QRS 161.22 (24.88) 162.51 (31.71) 167.89 (25.01) 162.32 (26.26)
[…]
The standard Kaplan-Meier procedures for describing survival probabilities do not provide proper results in the context of competing risks. In these settings, an alternative measure is the cumulative incidence function for each cause, defined as the cumulative probability of failure from a specific cause before a certain time. Figure 1 presents the cumulative incidence function for CV and N-CV death computed from the R package cmprsk (Gray, 2022). CV and N-CV death behave quite similarly, although cardiovascular death is always associated with a higher probability. It is important to note that after five years of the CRT intervention the difference between the two probabilities increases: between 0.15 and 0.20 for N-CV death and between 0.20-0.25 for CV death. From 7-8 years onwards there is an abrupt rise in CV death, while CV remains at around 0.2. The small number of patients we have from 10 years onwards does not produce consistent informative signals.
Figure 1: Estimation of the cumulative incidence function for time to CV (red) and non-CV (blue) death.
3 Bayesian competing risks modelling
3.1 Competing risk sampling model
Competing risk models are survival models in which the initiating event of the process contemplate different terminating events, which compete with each other to be the event that eventually occurs. There are different models suitable for analyzing this type of problems. The most popular, and the one we will use in our study, is the cause-specific hazard model. We assume $K$ terminating events or causes of failure.
We define $T_{k}$ as the random variable that describes the time to the terminating event $k$, and $\delta$ as the indicator variable for the event that has occurred. In particular, $\delta=k$ when the event $k$ has occurred and $\delta=0$ indicating the situation in which no event of interest has been recorded during the study period and, therefore, we have right censored information. Let $T=\mbox{min}\{T_{1},\ldots,T_{K}\}$. Cause-specific hazard modelling uses as a basic concept the joint distribution of $T$ and $\delta$ through the so-called cause-specific hazard function, which represents the hazard of failing from a given event in the company of all other competing events. The cause-specific hazard function for $T$ due to cause $k$ is defined as
$h_{k}(t\mid\bm{\theta}_{k})=\lim_{\varDelta t\to 0}\frac{P(t\leq T<t+\varDelta t,\delta=k\mid T\geq t,\,\bm{\theta}_{k})}{\varDelta t},$ (1)
where $\bm{\theta}_{k}$ is the subsequent parametric vector. It is important to remember that our inferential framework is Bayesian statistics. Consequently, the uncertainty associated with the parameters will be expressed in probabilistic terms and for this reason we will express in terms of conditional probability distributions all probabilistic elements, such as $h_{k}(t\mid\bm{\theta_{k}})$, that depend on parameters.
The most common modeling of the cause-specific hazard function for cause $k$ is through a proportional hazards model (Chen et al., 2014) as follows
$h_{k}(t\mid\bm{\theta}_{k})=h_{0,k}(t\mid\bm{\phi}_{k})\>\exp\{\bm{x}^{\top}_{k}\,\bm{\beta}_{k}\},$ (2)
where $h_{0,k}(t\mid\bm{\phi_{k}})$ is a baseline hazard function that depends on time and a parametric vector $\bm{\phi}_{k}$. It expresses the shape of the cause-specific hazard function. The vector $\bm{\beta}_{k}$ represents the regression coefficients associated to covariates $\bm{x}_{k}$ of covariates, and $\bm{\theta}_{k}=(\bm{\phi}_{k},\bm{\beta}_{k})$. Baseline hazard functions can be established through parametric or semiparametric approaches. Parametric models generate functions with rigid shapes that cannot admit irregular behavior. On the other hand, semi-parametric modeling provides very flexible functions but suffer from overfitting and instability (Ibrahim et al., 2014).
Outputs of interest in competing risks models include the overall survival function and the cumulative incidence function from each competing event $k$. The overall survival function for $T$ is
[…]
where now $\bm{\theta}=(\bm{\theta}_{1},\ldots,\bm{\theta}_{K})$.
The cumulative incidence function of $T$ due to cause $k$ describes the time-dependent probability that an individual experiences event $k$ in the presence of all the competing events as a function of time:
$F_{k}(t\mid\bm{\theta})=P(T\leq t,\>\delta=k\mid\bm{\theta})=\int_{0}^{t}h_{k}(u\mid\bm{\theta_{k}})\>S(u\mid\bm{\theta})\>\mbox{d}u.$
This function provides the probability associated with the occurrence of each of the competing events, $P(\delta=k\mid\bm{\theta})=\lim_{t\to\infty}F_{k}(t\mid\bm{\theta})$.
Competing risk models can also be seen as a particular multi-state model. In general, a multi-state process is a continuous time stochastic process $\{Z(t),t\geq 0\}$ with a finite state space $\Omega=\{0,1,\ldots,K\}$, where $Z(t)=k$ indicates that the process is in state $k$ at time $t$, or in terms of survival analysis that event $k$ has occurred at some time instant less than $t$. The initial distribution of the process is $P(Z(0)=1)=1$ and relevant transition probabilities between states are
$p_{(0,0)}(0,t\mid\bm{\theta})$ $=P(Z(t)=0\mid Z(0)=0,\bm{\theta})=P(T>t\mid\bm{\theta})=S(t\mid\bm{\theta}),$
$p_{(0,k)}(0,t\mid\bm{\theta})$ $=P(Z(t)=k\mid Z(0)=0,\bm{\theta})=P(T\leq t,\delta=k\mid\bm{\theta})=F_{k}(t\mid\bm{\theta}),\,\,k\neq 0.$
In many instances, for example in disease monitoring, it is often important for both the physician and the patient to assess the risks associated with disease progression not only from the beginning of the monitoring but also from subsequent periods. In these cases, the relevant transition probabilities are:
$p_{(0,0)}(s,t\mid\bm{\theta})=$ $\,\frac{P(Z(t)=0,Z(s)=0\mid\bm{\theta})}{P(Z(s)=0\mid\bm{\theta})}=\frac{S(t\mid\bm{\theta})}{S(s\mid\bm{\theta})},\,t\geq s,$
$p_{(0,k)}(s,t\mid\theta)=$ $\,\frac{P(Z(t)=k,Z(s)=0\mid\bm{\theta})}{P(Z(s)=0\mid\bm{\theta})}=\frac{P(s<T<t,\delta=k\mid\bm{\theta})}{S(s\mid\bm{\theta})}$
$=$ $\,\frac{F_{k}(t\mid\bm{\theta})-F_{k}(s\mid\bm{\theta})}{S(s\mid\bm{\theta})},\,t\geq s,\,k\neq 0.$
3.2 Prior and posterior distributions
Bayesian models include a sampling model and a prior distribution $\pi(\bm{\theta})$ for the unknown parameters that describes how much, little or no information is available about $\bm{\theta}$. We will work in a general scenario of prior independence between the parameters associated with the different competing risks as well as between the parameters governing the time-dependent structure of each hazard function and the regression coefficients of the variables, $\pi(\bm{\theta})=\prod_{k=1}^{K}\,\pi(\bm{\phi}_{k},\bm{\beta}_{k})=\prod_{k=1}^{K}\,\pi(\bm{\phi}_{k})\,\pi(\bm{\beta}_{k})$.
The database is built on the basis that survival times due to one of the causes provides censored information on the survival times of the other competing causes. The information provided by the data $\mathcal{D}$ is described in terms of the likelihood function of $\bm{\theta}$, $L(\bm{\theta})$, which is calculated through the hazard function (Chen et al., 2014). The prior distribution and the likelihood function are combined to obtain the posterior distribution of $\bm{\theta}$, $\pi(\bm{\theta}\mid\mathcal{D})$. This distribution is not analytical but an approximate sample $\{\bm{\theta}^{(1)},\ldots,\bm{\theta}^{(M)}\}$ of such a distribution can be obtained through Markov chain Monte Carlo, MCMC, methods. This distribution includes all available information and is the starting point for the analysis of the problem.
The overall survival function, cumulative incidence function, probability associated with the occurrence of each competing event as well as transitions probabilities presented before are relevant outputs that depend on $\bm{\theta}$. Consequently, we can use the posterior approximate sample from $\pi(\bm{\theta}\mid\mathcal{D})$ to generate approximate random samples for their subsequent posteriors, $\pi(S(t\mid\bm{\theta})\mid\mathcal{D})$, $\pi(F_{k}(t\mid\bm{\theta})\mid\mathcal{D})$, and $\pi(p_{(u,v)}(s,t\mid\bm{\theta})\mid\mathcal{D})$, respectively.
3.3 Model uncertainty
We assume uncertainty in the model, both in the specification of the baseline hazard function and in the set of covariates. In our problem, our main interest is in delving into the role of the covariates in the model and evaluating the real information that each of them provides about the survival times of interest. Model selection is a challenging problem in statistics from a methodological point of view and of great importance in modeling real problems. In current regression studies, increasingly larger sets of covariates are available in which covariates that really have a relevant impact on the variable of interest are mixed with covariates that barely provide information about the problem or whose information, although important, is redundant. Identifying the really valuable variables and discarding those that are not is a really necessary process. Regression models with many covariates are complex and rigid models that generate estimates with a lot of uncertainty due to situations of multicollinearity. They have a limited predictive ability.
Statistical literature is very rich in many and different proposals that address model uncertainty. We extend the proposal of García-Donato et al (2023) in Cox regression to competing risk models, and employ spike-and-slab methods for variable selection. Both procedures use a strategy whose first stage proposes a variable selection method and a subsequent analysis of the estimated model.
The approach proposed in García-Donato et al. (2023) is based on computing the posterior probability associated to each of the possible models (model space), from the null model containing no covariates to the full model with all covariates. The calculation of such posterior distribution uses a discrete uniform prior distribution over the model space and a multivariate normal distribution as the prior for the regression coefficients $\pi(\bm{\beta}_{k})=\mbox{N}(\bm{0},\,\Sigma),$ where $\Sigma^{-1}$ is expressed in terms of the precision matrix of the covariates. This variance-covariance matrix is based on the literature of conventional priors and the expected Fisher information matrix evaluated on the null model. Model search in the different $2^{p_{k}}$ models, where $p_{k}$ is the number of covariates in the cause-specific hazard for cause $k$, $k=1,\ldots,K$, is based on Gibbs sampling algorithm and the Rao-Blackwellized estimator (Gelfand and Smith, 1990; Ghosh and Clyde, 2011) is used for the estimation of the inclusion probability associated to each covariate.
The second selection method we will follow is based on spike and slab prior distributions for the set of regression coefficients. In this approach, the prior distribution associated to the $\bm{\beta}_{k}$ regression coefficients associated to cause $k$ is expressed as (Mitchell and Beauchamp, 1988)
$\pi(\bm{\beta}_{k}\mid\bm{\gamma})$ =$_{j=1}^{p_{k}}\,[(1-\gamma_{j})I_{0}(\beta_{kj})+\gamma_{j}\pi(\beta_{kj})],$
$\pi(\bm{\gamma}\mid\eta)$ =$_{j=1}^{p_{k}}\,\eta^{\gamma_{j}}(1-\eta)^{1-\gamma_{j}},$
$\eta$ $\sim\mbox{Be}(\alpha_{1},\alpha_{2}),$ (4)
where $I_{0}(\beta_{kj})$ is a mass point at zero that accompanies the spike, $\pi(\beta_{kj})$ is a wide density, and $\bm{\gamma}$ is a $\{0,1\}$ vector which defines the different $2^{p_{k}}$ models.
The two procedures we have used look at the covariates but are not designed, especially the first one, to assess the baseline hazard function. In this sense, as a complement to model selection we focused on other criteria to assess the performance of the model as a whole, with covariates and a baseline hazard function. The literature on this subject is extensive. We have selected the Deviance Information Criterion (DIC) (Spiegelhalter et al., 2002) and the Watanabe-Akaike Information Criterion (WAIC) (Watanabe, 2010). both based on information measures and the logarithm of the pseudomarginal likelihood (LPML) (Ibrahim et al., 2001) based on the predictive power of the model.
DIC measures the fit of the data to the model via the deviance with a penalty for the complexity of the model, the effective number of parameters. DIC has been a very popular procedure in the Bayesian world due to its practical relationship with MCMC procedures to approximate the posterior distribution and its good performance, basically in non-hierarchical scenarios as it is our case. Smaller values of DIC indicate a better-fitting model. WAIC criterion is an extension of the Akaike Information Criterion with a fully Bayesian conception and very good properties.
LPML is based on the Conditional Predictive Ordinate (CPO) associated to each observation (Ibrahim et al., 2001). It is defined as
$\mbox{LPML}=\sum_{i=1}^{n}\,\mbox{log}(\mbox{CPO}_{i}),$
where CPOi for observation $y_{i}$ in general univariate scenarios without censoring is defined in terms of the marginal posterior predictive distribution
$f(t\mid\mathcal{D}^{-(i)})=\int\,f(t\mid\bm{\theta})\,\pi(\bm{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\bm{\theta},$
where the $\mathcal{D}^{(-i)}$ denotes the data without the $i$th observation (leave-one-out procedure), and $f(t\mid\bm{\theta})$ is the subsequent probability sampling density. In the case of univariate survival times with right censoring, the CPO associated to observation $t_{i}$ is the value of that posterior marginal distribution evaluated at the observation $t_{i}$ if this time is observed and the posterior probability $P(T_{i}>t_{i}\mid\mathcal{D}^{(-i)})$ when $t_{i}$ is censored (Laud, 2014). Large CPOs values support the selected model because indicate a good tuning between the data and the model.
The above definition of CPO does not apply to competitive risk models in which there are different event causes and censoring. We propose an extension of the definition of the CPO associated to the $i$-th observation $t_{i}$ in a competing risk model as follows:
  • Individual $i$ has not experienced any of the events in the model during the period of the study. Consequently, observation $t_{i}$ for individual $i$ is censored for all competing events ($\delta_{i}=0$). Its associated CPO is the value at $t_{i}$ of the posterior predictive distribution
    $P(T>t\mid\mathcal{D}^{-(i)})=S(t\mid\mathcal{D}^{-(i)})=\int S(t\mid\bm{\theta})\,\pi(\bm{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\bm{\theta},$
    where $S(t\mid\bm{\theta})$ is the overall survival function in(3).
  • Individual $i$ has experienced the event $k$ in time $t_{i}$ within the period of the study. In this case the CPO associated to this observation is the value at $t_{i}$ of the posterior predictive subdensity function
    $\ f_{k}(t\mid\mathcal{D}^{-(i)})=\int f_{k}(t\mid\bm{\theta})\,\pi(\bm{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\bm{\theta}=\int h_{k}(t\mid\bm{\theta})\,S(t\mid\bm{\theta})\,\pi(\bm{\theta}\mid\mathcal{D}^{-(i)})\,\mbox{d}\bm{\theta},$
    where $f_{k}(t\mid\bm{\theta})=h_{k}(t\mid\bm{\theta})\,S(t\mid\bm{\theta})$ is known as the subdensity sampling function for time-to-event $k$, defined in terms of the cause-specific hazard function for cause $k$ in (1) and the overall survival function (3).
CPOs and LPML ca be used to compare the predictive performance of different models as well as to search for outliers. We will use an approximation for the calculation of the CPOs based on the MCMC outputs (Ntzoufras, 2008).
4 Bayesian competing risk analysis in action: Cardiovascular and non-cardiovascular death
4.1 Model specification
We return to the cardiac resynchronization therapy study, whose data we will analyse using a Bayesian competing model with two competing events, CV and N-CV death. Each patient undergoing CRT (initial state) can access any of them after having surgery. We define $T_{C}$ and $T_{NC}$ as the survival time associated with CV and N-CV death, respectively. We express the cause-specific hazard function for CV and N-CV death in terms of a Cox proportionals hazards modelling,
$h_{c}(t\mid\bm{\phi}_{c},\,\bm{\beta}_{c})=\,h_{0,c}(t\mid\bm{\phi}_{c})\,\exp\{\bm{x}^{\top}_{c}\bm{\beta}_{c}\},$
$h_{nc}(t\mid\bm{\phi}_{nc},\,\bm{\beta}_{nc})=\,h_{0,nc}(t\mid\bm{\phi}_{nc})\,\exp\{\bm{x}^{\top}_{nc}\,\bm{\beta}_{nc}\},$
where $h_{0,c}(t\mid\phi_{c})$ ($h_{0,nc}(t\mid\phi_{nc})$) is the baseline hazard function for cardiovascular (non-cardiovascular) death and the element with covariates is :
$\bm{x}^{\top}_{c}\bm{\beta}_{c}=\,\,\beta_{c.1}I_{Man}+\beta_{c.2}I_{Pace}+\beta_{c.3}I_{Ische}+\beta_{c.4}I_{III}+\beta_{c.5}I_{IV}$
$\hskip 24.18501pt+\,\,\beta_{c.6}\,\mbox{Age}+\beta_{c.7}\,\mbox{Crea}+\beta_{c.8}\,\mbox{LBBB}+\beta_{c.9}\,\mbox{PAF}++\beta_{c.10}\,\mbox{QRS},\vspace*{0.5cm}$
$\bm{x}^{\top}_{nc}\,\bm{\beta}_{nc}=\,\,\beta_{nc.1}I_{Man}+\beta_{nc.2}I_{Pace}+\beta_{nc.3}I_{Ische}+\beta_{nc.4}I_{III}+\beta_{nc.5}I_{IV}+$
$\hskip 32.72049pt+\,\,\beta_{nc.6}\,\mbox{Age}+\beta_{nc.7}\,\mbox{Crea}+\beta_{nc.8}\,\mbox{LBBB}+\beta_{nc.9}\,\mbox{PAF}++\beta_{nc.10}\,\mbox{QRS}.$
The term $I_{A}$ is an indicator function whose value is 1 if the subsequent individual has the characteristic $A$ and 0 otherwise. In particular, the five indicator functions in the two expressions above, $I_{Man}$, $I_{Pace}$, $I_{Ische}$, $I_{III}$, and $I_{IV}$ refer to being man, having a pacemaker implanted, ischaemic heart disease, and functional classification III and IV, respectively. Baseline covariates Age, Crea, LBBB, PAF, and QRS describe for each individual their age, level of creatinite, LBBB, previous AF, and QRS, respectively. The reference category for both causes of death is being woman, having an implanted ICD, dilated heart disease, and functional classification II.
Under the frequentist statistics it is not usual to specify the baseline hazard function. This is not the case for Bayesian inference, which requires the choice of a baseline hazard function to perform a full Bayesian analysis of the problem. In this regard, we account for two different specifications of the baseline hazard function associated to each competing event: a Weibull distribution and a piecewise function. In the case of a Weibull distribution with shape $\alpha_{c}$ and $\alpha_{nc}$ and scale parameter $\lambda_{c}$ and $\lambda_{nc}$ for the CV and N-CV death, respectively, the baseline hazard function for $T_{C}$ and $T_{NC}$ is
$h_{0.c}(t\mid\bm{\phi}_{c})=\lambda_{c}\,\alpha_{c}\,t^{\alpha_{c}-1},$
$h_{0.nc}(t\mid\phi_{nc})=\lambda_{nc}\,\alpha_{nc}\,t^{\alpha_{nc}-1},\,\,\,t>0.$
This baseline hazard function is a traditional model for survival data. It is very simple and easy to interpret but lacks flexibility to represent non-monotonic risks.
Piecewise functions are defined through polynomial functions. They provide flexible modeling that can include monotonic and non-monotonic behaviors. It is one of the traditional alternatives to the Weibull modelling used in Bayesian inference. We use the piecewiese constant function for both survival times which we define as:
$h_{0.c}(t\mid\bm{\phi}_{c})=\mbox{$\sum$}_{r=1}^{R_{c}}\,\gamma_{c.r}\,I_{{(a_{c.r-1},a_{c.r}]}},$
$h_{0.nc}(t\mid\bm{\phi}_{nc})=\mbox{$\sum$}_{r=1}^{R_{nc}}\,\gamma_{nc.r}\,I_{{(a_{nc.r-1},a_{nc.r}]}},$ (5)
where $0=a_{c.0}\leq a_{c.1}\leq\ldots\leq a_{c.R_{c}}$ and $0=a_{nc.0}\leq a_{nc.1}\leq\ldots\leq a_{nc.R_{nc}}$ is a partition of the relevant time axis for CV and N-CV death, respectively. In this case, the indicator function $I_{A}$ is 1 if $t$ belongs to interval $A$ and zero otherwise. In our study, we have considered partitions determined by knots 0, 3, 6, 12, 24, 36, 48, and 60 month for CV death and knots 0, 24 and 48 months for N-CV death. They have been selected by the cardiologist who collaborates in the project who is also a co-author of this paper.
Table 2: Approximate posterior distribution for the full model parameters, considering a Weibull and a piecewise function for the baseline hazard function.
Weibull baseline Piecewise baseline
Covariates Mean SD Mean SD
Cardiovascular death
Weibull shape 0.785 0.091
Weibull scale 0.009 0.006
Age 0.301 0.176 0.268 0.174
Sex (Man) 0.322 0.457 0.221 0.452
Creatinine 0.493 0.126 0.431 0.121
CRT Device (Pacemaker) 0.663 0.311 0.662 0.304
Etiology (Ischemic) 0.585 0.311 0.590 0.302
LBBB -0.019 0.291 -0.101 0.295
NYHA (Class III) 1.221 0.639 1.119 0.642
NYHA (Class IV) 1.734 0.755 1.793 0.750
Previous AF 0.315 0.290 0.309 0.290
QRS 0.008 0.139 -0.001 0.137
Non-Cardiovascular death
Weibull shape 1.234 0.166
Weibull scale 0.007 0.005
Age 0.688 0.242 0.577 0.233
Sex (Man) 0.092 0.427 0.047 0.422
Creatinine 0.306 0.242 0.120 0.226
CRT Device (Pacemaker) -0.096 0.364 -0.212 0.360
Etiology (Ischemic) 0.387 0.381 0.415 0.377
LBBB 0.518 0.369 0.570 0.369
NYHA (Class III) 0.527 0.589 0.660 0.572
NYHA (Class IV) -1.316 1.406 -0.997 1.400
Previous AF 0.556 0.367 0.623 0.357
QRS 0.199 0.170 0.179 0.163
We need to complete the Bayesian model with the specification of a prior distribution for all the parameters of the model, $\bm{\theta}=(\bm{\phi}_{c},\bm{\phi}_{nc},\bm{\beta}_{c},\bm{\beta}_{nc})$. We assume prior independence and have selected, as much as possible, non-informative prior distributions that are standard in this application environment in order to give all the prominence of the inferential process to the data. We have elicited normal distributions $N(0,0.001)$ for the coefficients of regression in $\bm{\beta}_{c}$, $\bm{\beta}_{nc}$. In the case of Weibull baseline risk functions, we have elicited gamma distributions $\text{Ga}(0.01,0.01)$ for the $\lambda$’s and uniforms $\text{U}(0,10)$ for the $\alpha$’s. Alternatively, in the case of the risk functions defined through the piecewise functions in (5), we have selected gamma distributions $\text{Ga}(0.01,0.01)$ for the $\phi$’s coefficients. We are working with two models that are identical except for the specification of the baseline risk function. The posterior distribution $\pi(\bm{\theta}\mid\mathcal{D})$ of each of them has been approximated via Bayesian inference through MCMC with JAGS software (Plummer, 2003) through the rjags package (Plummer, 2023) in the R version 4.1 (R Core Team, 2023). Three MCMC chains, 100,000 iterations, 1,000 burning iterations and a thinning interval of 10 have been used to compute each model (Álvares et al., 2021). For each model, convergence has been checked via the Gelman-Rubin statistic.
Table LABEL:tab:2 includes a summary of the approximate posterior distribution of the parameters of both models (whose only formal difference is in the specification of the hazard function). In particular, the posterior mean and standard deviation of the coefficients of both models are presented. Regarding the results, we would first like to comment on the similarity between the estimates of the parameters associated with the different covariates in both models, both for CV and N.CV death. It is also interesting to note that in both types of death, the posterior standard deviation associated with the categorical covariates is greater than that estimated for the quantitative covariates and that greater uncertainty is observed in the Cr death group. In the case of CV death, the categories of being a man, have a pacemaker device, an ischaemic etiology of the disease, and a functional classification III and IV generate an increase in the hazard function and, therefore, lower survival times. Quantitative covariates also have a negative impact on survival, except for LBBB and certainly QRS. In the N-CV death group, it is worth noting the low relevance of gender and the type of device implanted. In addition, note that having a diagnosis in functional classification IV decreases the risk of death, but age increases this risk notably.
Some alternative prior distributions on model parameters have been tested in order to check the stability of the obtained posterior distribution: $N(0,0.01)$ and a $U(-10,10)$ for the regression coefficients. Regarding the coefficients of the piecewise function, two additional prior specifications have been considered as described in Lázaro et al. (2020):
  1. 1.
    $\pi(\gamma_{c.r})=\mbox{Ga}(\omega_{c.0}\,\eta_{c.0}(a_{c.r}-a_{c.r-1}),\,\omega_{c.0}(a_{c.r}-a_{c.r-1}))$ and $\pi(\gamma_{nc.r})=\mbox{Ga}(\omega_{nc.0}\,\eta_{nc.0}(a_{nc.r}-a_{nc.r-1}),\,\omega_{nc.0}(a_{nc.r}-a_{nc.r-1}))$, where $\omega_{c.0}=\omega_{nc.0}=0.01$, $\eta_{c.0}=0.69315/\tilde{t}_{c}$, and $\eta_{nc.0}=0.69315/\tilde{t}_{nc}$ being $\tilde{t}_{c}$, and $\tilde{t}_{nc}$ the median survival time for the cardiovascular and non-cardiovascular cause, 1.79 and 2.38 years, respectively.
  2. 2.
    $\pi(\gamma_{c.r}\mid\gamma_{c.1},\,\gamma_{c.2},\ldots,\gamma_{c.r-1})=\mbox{Ga}(\eta_{c.0},\eta_{c.0}/\gamma_{c.r-1})$ and $\pi(\gamma_{nc.r}\mid\gamma_{nc.1},\,\gamma_{nc.2},\ldots,\gamma_{nc.r-1})=\mbox{Ga}(\eta_{nc.0},\,\eta_{nc.0}/\gamma_{nc.r-1})$, where $\eta_{c.0}=\eta_{nc.0}=0.01$ and $\pi(\gamma_{c.1})=\pi(\gamma_{nc.1})=\text{Ga}(0.01,0.01)$.
We have not detected any relevant change in the posterior distribution obtained from these new prior distributions compared with the results presented in Table LABEL:tab:3. We therefore have results that we could describe as robust.
4.2 Model uncertainty
Figure 2: Estimated inclusion probabilities of each of the covariates in the model using the Bayes factors based procedure in relation to the number of Gibbs iterations. Both competing events, (A) cardiovascular and (B) non-cardiovascular death, are considered.
As we have introduced in Section 3, we have used two variable selection procedures: the proposal of Garcia-Donato et al. (2023) through the Bayes factor and the spike-and-slab method. Regarding the first proposal, it should be remembered that it does not consider the baseline hazard function and only takes into account the covariates. A total number of 1000 Gibbs iterations and a burn-in period of 300 were used. The full model contained 10 covariates and the initial model dimension was set as 5. The stability of the inclusion probabilities was approximately achieved on the first 300 iterations as seen in Figure 2. The high probability of inclusion of Creatinine and Age in the case of CV and N-CV, respectively, can be observed. The probability of inclusion of the remaining covariates is very low for N-CV death. When it comes to CV death we can observe inclusion probabilities around 0.5 for CRT Device, Etiology and Age and very small for the remaining covariates.
Table LABEL:tab:3 shows the five competing risk models that the García-Donato et al. (2023) procedure selects with the highest probability. We again confirm that the results obtained for both causes of death are very different. We have little uncertainty about the best model in the case of N-CV death, with only covariate Age; and a little more uncertainty in the case of CV death: Creatinine appears in all the models and, in fact, the model with the highest probability only includes this covariate. CRT Device and Etiology appear in three of the five models and Age in only one.
Table 3: Posterior model probabilities of the five most probable competing risk models
Model Post. probability
Cardiovascular death
Creatinine 0.1913
Creatinine + Etiology 0.1452
Creatinine + CRT Device + Age + Etiology 0.1113
Creatinine + CRT Device 0.0815
Creatinine + CRT Device + Etiology 0.0465
Non-cardiovascular death
Age 0.6075
Age + NYHA(IV) 0.0790
Age + NYHA(III) 0.0592
Age + Etiology 0.0493
Age + Sex 0.0423
The spike and slab method applies to all elements in the model, the baseline risk function and the covariates. We recall that we have worked with two different formulations of the basal risk function, one through the Weibull distribution and the second through a piecewise-defined function. A $N(0,0.001)$ prior distribution has been selected for modeling each slab term and the beta distribution Beta(1,1) has been elicited associated to probability $\eta$ in (4).
Table 4: Inclusion probabilities of covariates associated with cardiovascular and non-cardiovascular death determined through the Bayes factor and the spike and slab method, this latter with a Weibull and a piecewise baseline risk function.
Spike-and-slab prior
Bayes Factor Weibull Piecewise
CV death N-CV death CV death N-CV death CV death N-CV death
Age 0.412 0.947 0.021 0.984 0.037 0.990
Sex 0.216 0.077 0.048 0.012 0.055 0.022
Creatinine 0.963 0.087 0.998 0.011 0.993 0.017
CRT Device 0.507 0.059 0.095 0.010 0.121 0.008
Etiology 0.447 0.091 0.190 0.020 0.162 0.015
LBBB 0.099 0.082 0.012 0.014 0.009 0.007
NYHA (Class III) 0.225 0.102 0.033 0.022 0.055 0.071
NYHA (Class IV) 0.214 0.133 0.028 0.051 0.031 0.082
Previous AF 0.176 0.109 0.028 0.028 0.021 0.053
QRS 0.100 0.086 0.004 0.010 0.010 0.003
Table LABEL:tab:4 shows the estimated probability of inclusion in the model of each of the covariates provided by the Bayes factor method and by a spike distribution method, the latter with the Weibull and the piecewise baseline risk function. We can observe that the results obtained with the spike are much more drastic than with the procedure based on the Bayes factor because they only give a high inclusion probability estimate to the covariate Creatinine in the case of CV death and to Age in the case of N-CV death. Both covariates are also the favorites of the Bayes factor. No relevant differences were observed between the two modeling approaches based on the spike-and-slab method. The model with the highest posterior probability and the model that includes all covariates with inclusion probability 0.5 or higher (Median posterior probability model, Barbieri and Berger, 2004) coincide in this case: Creatinine for CV death and Age for N-CV death.
All the information obtained on previous analyses matches with clinical knowledge. According to the clinicans criteria, the best model would be considering Age, Creatinine, CRT Device, and Etiology, and adding the covariate NYHA functional class for CV death due to their clinic relevance and leaving the proposal of both methods in the case of N-CV death with only Age.
As a final step, we used the DIC, WAIC and LPML criteria to compare four models: the full model with all covariates, the model that only includes covariates that have a probability of inclusion greater than 0.5, the model with covariates with inclusion probabilities greater than 0.4, and the previous model with the NYHA covariate added. In all models, we have considered a Weibull and a piecewise polynomial baseline risk function. All the models have a fairly similar behavior with the three criteria, although the models with Weibull baseline risk function are always better and also, in all cases, the last one is the best. Finally, we decided to select the latter model: Weibull baseline hazard function for all events and covariates. Age, Creatinine, Device, Etiology and NYHA for CV death, and Age for N-CV death.
4.3 Selected model and posterior outputs
Table LABEL:tab:5 shows a summary of the approximated posterior distribution for the parameters of the selected model. We have used JAGs in the same way as with the full model to obtain an approximate sample from the subsequent posterior distribution. If we observe the results of the complete model and the selected model, we can see stability and similarity in the information corresponding to the common covariates in both events, especially for CV death. In this case, the ischemic etiology is associated with an increased risk of death compared to the dilated etiology. This also occurs with the CRT pacemaker with regard to the defibrillator device. Age, Creatinine and NYHA functional classification III and mainly IV have also a positive impact on CV death. Age is the only relevant covariate for N-CV death, with a strong positive impact on the risk of death.
Table 5: Summary of the approximate posterior distribution for the selected model.
Mean SD Median CI 95%
Cardiovascular death
Weibull shape 0.780 0.090 0.777 [0.611, 0.965]
Weibull scale 0.012 0.007 0.010 [0.002, 0.030]
Age 0.326 0.172 0.323 [-0.004, 0.674]
Creatinine 0.505 0.120 0.505 [0.267, 0.736]
CRT Device (Pacemaker) 0.689 0.316 0.687 [0.087, 1.315]
Etiology (Ischemic) 0.616 0.287 0.616 [0.055, 1.187]
NYHA (Class III) 1.284 0.644 1.232 [0.177, 2.712]
NYHA (Class IV) 1.753 0.769 1.718 [0.345, 3.375]
Non-Cardiovascular death
Weibull shape 1.218 0.158 1.214 [0.920, 1.537]
Weibull scale 0.022 0.007 0.021 [0.011, 0.039]
Age 0.826 0.224 0.819 [0.403, 1.283]
Figure 3: Posterior mean and 95% credible intervals of transition probabilities to CV death (red), to N-CV death (blue) and survival (green) for two different patient profiles. (A) 65 year-old patient with serum creatinine of 1.2 mg/dL, ischemic etiology, defibrillator implanted and NYHA III; (B) 75 year-old patient with serum creatinine of 1.5 mg/dL, ischemic etiology, pacemaker implanted and NYHA IV.
Figure 3 shows the posterior mean for the relevant transition probabilities of the model: from the initial state defined by the time in which the CRT implant is performed until CV, N-CV death or the permanence in that initial state in which the patient continues with life. We have computed these probabilities for two different patient profiles. On the one hand, 65 year-old patients, with ischemic disease etiology, pacemaker implanted, NYHA class III and a low baseline creatinine level (1.2 mg/dL ). In the second group of patients we included 75 year-old patients, ischemic etiology, pacemaker, NYHA class IV and a high creatinine level (1.5 mg/dL). From now on, we will call the patients in the first group type A patients and the second group type B patients.
The behavior of transition probabilities in both types of patient profiles is very different. Patients with a type A profile have better prognoses than those with a B profile. The probability of staying alive after CRT implantation in patients with profile A is greater than 0.5 until approximately 13 years after implantation, whereas for patients with profile B it is only greater than 0.5 until approximately 3 years after implantation. The probability of being alive 5, 10 and 15 years after implantation is around 0.76, 0.57, and 0.46 in type A patients, and around 0.31, 0.13, and 0.05 in type B patients, respectively. These probabilities are also known in clinical terms as 5, 10 and 15-year survival. The probability of CV death before 5, 10 or 15 years after implantation is around 0.15, 0.25 and 0.29 in type A patients. These probabilities become 0.57, 0.67 and 0.73 in type B patients. As for N-CV death, in type A patients there is an estimated probability of 0.09, 0.18 and 0.23 that death will occur in the first 5, 10 and 15 years after the implant. These estimates are 0.12, 0.21 and 0.22 in type B patients. There do not seem to be many differences in the evolution of N-CV death in both patient profiles. The largest differences are found in the survival and CV death probabilities.
Figure 4 includes, for patients type A and B, the posterior mean of the conditional transition probabilities to a CV or a N-CV death as well as the permanence as a living individual in a certain period of time given that he/she has been implanted and has survived a known length of time after implantation. The term conditional here is justified by the assumption that the patient continues to live for a certain period of time after implantation, in our case 3, 5 and 8 years. If we look at type A patients, we recall that their 10 and 15-year survival was around 0.57 and 0.46, respectively. If after 3, 5 or 8 years a type A patient had survived, the expected posterior probability that he would still be alive 10 years after implantation increases to around 0.70, 0.76 and 0.90, respectively. These expected probabilities for a 15-years period of time are about 0.55, 0.60 and 0.73, respectively.
Figure 4: Posterior mean and 95% credible intervals of conditional transition probabilities to cardiovascular death (red line), to non-cardiovascular death (blue line) and survival (green line) for two different patient profiles. 3, 5 and 8 years are considered as the initial survival time. (A) 65 year-old patient with serum creatinine of 1.2 mg/dL, ischemic etiology, defibrillator implanted and NYHA III; (B) 75 year-old patient with serum creatinine of 1.5 mg/dL, ischemic etiology, pacemaker implanted and NYHA IV.
In the case of CV death the survival of type A patients was less than 10 and 15 years with probabilities 0.24 and 0.29. These probabilities turn out to be 0.67 and 0.73 for type B patients. Adding the condition that the patient is still alive 3 years after implantation, these probabilities become 0.15 and 0.23 for type A patients and 0.54 and 0.62 for type B patients,. respectively. In the case of having lived more than 5 years after transplantation, the probabilities of living less than 10 years become 0.12 and 0.17 for type A patients, and 0.46 and 0.61 for type B patients. Finally, for patients who survive more than 8 years after implantation, the probability of a CV-death less than 10 or 15 years after implantation is 0.05 and 0.13 for type A patients and 0.24 and 0.50 for type B patients, respectively.
The conditional probabilities associated with N-CV death are the most stable. We know that in type A patients the probability of N-CV death before 10 to 15 years after implantation is around 0.19 and 0.25, and approximately 0.21 and 0.23 in type B patients. If we add the condition of having survived more than 3, 5 and 8 years after implantation these probabilities become 0.15 and 0.23, 0.12 and 0.23, and 0.05 and 0.14 in type A patients, respectively. In patients type B, the aforementioned probabilities are approximately 0.23 and 0.26, 0.21 and 0.26, and 0.11 and 0.25, respectively.
5 Conclusions
We have proposed a Bayesian approach to competing risk survival models through the so-called cause-specific hazard models which we have complemented through a multistate model perspective. We have carried out an inferential process with non-informative prior distributions and MCMC methods to approximate the respective posterior distribution. Outputs of interest in the statistical study also include the posterior distribution for the overall survival function of the time in which the first event occurs and the cumulative incidence function of that survival time due to each event-cause. In addition, we have discussed the posterior distribution for the transition probabilities between the different states of the multistate model both from the initial event and the case where the individual under consideration remains under the initial situation for a certain period of time.
Model uncertainty has been adressed for both variable selection and the specification of the baseline hazard function in the subsequents proportional hazard functions associated to the time in wchich the event occurs for each cause. The proposal by García-Donato et al (2023) with regard Bayes factor for univariate Cox regression is generalised to competing risk models. An extension of the CPO, which is the basis for logarithm of the pseudomarginal likelihood criterion based on the predictive power of the model, for competing risks scenarios is also presented.
We have applied all the presented procedures to data from a retrospective follow-up study conducted in a cardiac institution in Galicia (Spain) between January 2005 and April 2015 to assess time to CV and N-CV death in patients undergoing CRT. The results obtained have been very positively assessed by the cardiologist of the team insofar as they could be very useful in clinical practice. It is worth highlighting the selection of the set of covariates chosen in the survival times associated with both causes of death and the information provided by the posterior distribution related to the relevant transition probabilities for different patient profiles.
{acknowledgement}
This research was supported by the Xunta de Galicia via an industrial doctoral grant to the first author, grants SBPLY/17/180501/000491 and SBPLY/21/180501/000241, funded by Consejería de Educación, Cultura y Deportes (JCCM, Spain) and Fondo Europeo de Desarrollo Regional, and the Ministerio de Ciencia, Innovación y Universidades through the project “PID2022-136455NB-I00 MCIN/AEI/10.13039/501100011033/FEDER, UE” and the European Regional Development Fund.
Conflict of Interest
The authors have declared no conflict of interest.
Data availability
The JAGS codes for reproducing the analyses are available on GitHub: https://github.com/jegubo/BUGS_Competing.git.